OpenStudy (anonymous):

so implicit differentiation again...xy+y^2=4

OpenStudy (astrophysics):

Tis astro again to save the day

OpenStudy (astrophysics):

You will need to apply the product rule and note \[(y^2)' = 2y \frac{ dy }{ dx }\]

OpenStudy (astrophysics):

\[y' = \frac{ dy }{ dx }\]